# Behavioral Refinement via Interpolant-based Policy Diffusion

### v1

@@ Abstract
Imitation learning empowers artificial agents to mimic behavior by learning from demonstrations. Recently, diffusion models, which have the ability to model high-dimensional and multimodal distributions, have shown impressive performance on imitation learning tasks. These models learn to shape a policy by diffusing actions (or states) from standard Gaussian noise. However, the target policy to be learned is often significantly different from Gaussian and this mismatch can result in poor performance when using a small number of diffusion steps (to improve inference speed) and under limited data. The key idea in this work is that initiating from a more informative source than Gaussian enables diffusion methods to mitigate the above limitations. We contribute both theoretical results, a new method, and empirical findings that show the benefits of using an informative source policy. Our method, which we call BRIDGeR, leverages the stochastic interpolants framework to bridge arbitrary policies, thus enabling a flexible approach towards imitation learning. It generalizes prior work in that standard Gaussians can still be applied, but other source policies can be used if available. In experiments on challenging benchmarks, BRIDGeR outperforms state-of-the-art diffusion policies and we provide further analysis on design considerations when applying BRIDGeR.
@@ I Introduction
An examination of existing diffusion-style imitation learning reveals a fundamental issue: these models learn to shape a policy starting from standard Gaussian noise, which is often starkly different from the intended policy or action distribution. The key insight in our work is that initiating from Gaussian noise isn’t a prerequisite. To explore this, we move beyond the conventional diffusion framework and employ stochastic interpolants [2] for bridging arbitrary densities within finite time (Fig. 1). This approach allows us to leverage stochastic source policies, enabling the diffusion process to begin from a more informative starting point. We find that this shift retains the inherent advantages of diffusion-style imitation learning, but positively impacts inference time and performance. Our approach also generalizes prior work in diffusion-based imitation learning since if no source policy is available, simple distributions such as the Gaussian can be used.
@@
In this paper, we first contribute a theoretical analysis of the impact of different source policies in diffusion. In brief, we find that under reasonable assumptions, selecting a better source policy results in better target policies. We then turn to a practical approach for incorporating source policies into diffusion methods. Applying the stochastic interpolants framework [2] to imitation learning, we derive a new method called BRIDGeR (Behavioral Refinement through Interpolant-based Diffusion for Generative Robotics). To our knowledge, our work is the first application of this bridging methodology to imitation learning, contrasting with its previous use in simple synthetic tasks [58] and image generation [58, 24]. In addition to standard neural architecture design for the learnt forward model, the stochastic interpolant framework relies on several critical design choices including the source policy and interpolant. The interpolant dictates how a sample transitions from the source to the target distribution, with the transition modulated by noise introduced through time-dependent Gaussian latent variables [2].
@@
We contribute a systematic empirical study of the effects of using source policies (and other design elements) on a diverse set of robot tasks, including the Franka kitchen benchmark, grasp generation, and challenging tasks using a robot hand. Overall, the experimental results coincide with our theoretical findings; Gaussians were seldom the most effective source distribution and surprisingly, even simple heuristic distributions resulted in superior learnt policies compared to the Gaussian. We demonstrate that given a good source policy, BRIDGeR surpasses existing state-of-the-art diffusion policies. Additionally, we discuss the effects of the interpolant function when learning highly multi-modal behaviors.
@@
II Preliminaries: Background and Related Work
@@ II-A Problem Formulation
In imitation learning, we wish to learn a policy from expert demonstrations. Let $\pi_{1}(a|x)$ be an expert policy; it captures the probability of an expert selecting an action $a$ given an observation $x$. Suppose we have a dataset $\mathcal{D}=\{\smash{x^{(i)}},\smash{a^{(i)}}\}_{i=1}^{N}$ drawn from $\pi_{1}(a|x)$ and let $\pi_{0}(a|x)$ be a distribution that we can easily sample $a$ from.
@@ IV-A Stochastic Interpolants for Imitation Learning
Let $(C^{r}(\mathbb{R}^{n}))^{m}$ be the space of $r$ continuously differentiable functions from $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$ and $(C_{0}^{r}(\mathbb{R}^{n}))^{m}$ as the space of compactly supported and $r$ continuously differentiable functions from $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$. We extend the definition of stochastic interpolant in [2] to condition on observation $x$, similar to the concurrent work [18].
@@
     $L_{b}(\theta)=\frac{1}{N}\sum\left(b_{\theta}(t,a_{t},x)-(\partial_{t}I+\dot{\gamma}z)\right)$.
     $L_{s}(\eta)=\frac{1}{N}\sum\left(s_{\eta}(t,a_{t},x)+\gamma^{-1}z\right)$.
     $L_{v}(\phi)=\frac{1}{N}\sum\left(v_{\phi}(t,a_{t},x)-\partial_{t}I\right)$.
@@ IV-B Design Decisions
V Related Work
BRIDGeR builds upon a large body of work in imitation learning [19, 57, 49], specifically behavior cloning.
Behavior cloning has had a long history, with early methods adopting supervised learning techniques, principally regression [4, 33, 33, 56, 12, 37, 46]. Although efficient, regression-based methods were generally unable to capture multi-modal behavior. Later methods used classification methods, by discretizing the action space [40, 41, 3, 50], to address this limitation but were sensitive to the hyperparameters (such as the level of discretization) and struggled with high-precision tasks [11].
Modern imitation methods employ generative models, starting with Gaussian Mixture Models [27], then transitioning to more powerful Energy-Based Models (EBMs) [7, 11, 20, 43] and Diffusion Models [39, 6, 31, 30]. Compared to EBMs, diffusion models have demonstrated better training stability [6] and effectiveness in generating consistent, multi-modal action sequences for visual-motor control [39, 6, 31, 30] and hierarchical, long-horizon planning tasks [35, 15, 29, 5, 51].
Recent progress has been made in reducing the computational costs of diffusion methods by adapting the number of diffusion steps [17] or leveraging DDIM [36].
Unlike the above methods, BRIDGeR is the first to generalize diffusion-type policy learning to exploit source stochastic policies. As we will see, this leads to better performance compared to strong baselines (such as DDIM) when informative policies are available, especially when using a small number of diffusion steps.
Since BRIDGeR adapts existing distributions, it can be seen as a few-shot learner, specifically for imitation learning [9, 49]. Existing work on few-shot imitation learning requires either specific prior policies [10, 55, 23] or a hierarchical structure [52]. In contrast, BRIDGeR only requires that we are able to sample from the source policy. For fair comparison against few-shot methods, we adopt residual learning which is used for few-shot reinforcement learning [42, 1, 21], planning [28], and shared autonomy [38, 54].
VI Experiments
This section describes experiments designed to evaluate the performance of BRIDGeR relative to recent methods, particularly diffusion-based imitation learning. More importantly, we aimed to test our hypothesis that leveraging better source distributions within a diffusion framework can lead to better policies. We further hypothesized that BRIDGeR performs better than DDIM policies given a small number of diffusion steps, and limited data. Finally, we sought to examine the effect of decision decisions, particularly the interpolant choice.
In the following, we first give an overview of our experimental setup (details relegated to Appendix) followed by our main results.
VI-A Domains
@@
Table I: Average task performance of BRIDGeR against current state-of-the-art methods under a different number of diffusion steps when trained with the Large dataset. Overall, BRIDGeR outperforms all competing methods. Results are similar for the small and medium datasets with complete results in the Appendix.
@@
Table II: Success rate (averaged over $100$ grasps on ten test objects). BRIDGeR significantly outperforms DDIM and Residual Policy across the number of diffusion steps. Compared to SE3, BRIDGeR achieve higher success rate when the number of diffusion steps is small. We show up to $k=160$ steps to be consistent with prior reported results [47].
@@
VI-D Main Results and Discussion
Figure 7: Fifteen sampled grasps across the diffusion steps for BRIDGeR, DDIM and SE3. We visualize $15$ grasps samples every $2$ diffuse steps until $k=10$.
@@ VI-D Main Results and Discussion
Figure 9: Average Earth Mover’s Distance between generated grasps pose and target grasp poses on Seen Categories in Grasp Generation task ($k=160$ for SE3, DDIM and BRIDGeR). Lower EMD indicates BRIDGeR is closer to the dataset distribution.
Figure 10: Task performance of BRIDGeR starting from different source policies using the Power3 interpolant. In general, with better source distribution, BRIDGeR achieves higher task performance.
@@ VII Conclusions and Future Work
In this work, we investigate the potential of integrating informative source distributions into diffusion-style imitation learning. We provide theoretical results that support this idea and propose BRIDGeR, a stochastic interpolant method for imitation learning. Our experiments results show that BRIDGeR outperforms strong baselines, including state-of-the-art diffusion policies on various benchmark tasks. We provide additional analyses on our experimental results to elucidate the effect of various design decisions within BRIDGeR.
Limitations and Future Work. Here, we have shown that leveraging prior knowledge (in the form of source policies) improves learned diffusion policy performance. Here, the prior knowledge can be obtained either from human experts or existing datasets. It would be interesting to explore other forms of prior knowledge, e.g., policies designed for other tasks and those constructed using foundation models. Specific to BRIDGeR, we explored a salient but limited set of design considerations; future work can examine more elaborate interpolant functions, along with more in-depth analysis of noise schedules and diffusion coefficients. Moreover, our evaluations comprise simulation benchmarks as our study involved tens of thousands of trials; recent work [6] has shown that diffusion policies can work on real robots and we plan to conduct real world trials using BRIDGeR in subsequent work. Finally, we plan to extend BRIDGeR to enable classifier-guided sampling.

### v2

# Don’t Start from Scratch: Behavioral Refinement via Interpolant-based Policy Diffusion

@@ Abstract
Imitation learning empowers artificial agents to mimic behavior by learning from demonstrations. Recently, diffusion models, which have the ability to model high-dimensional and multimodal distributions, have shown impressive performance on imitation learning tasks. These models learn to shape a policy by diffusing actions (or states) from standard Gaussian noise. However, the target policy to be learned is often significantly different from Gaussian and this mismatch can result in poor performance when using a small number of diffusion steps (to improve inference speed) and under limited data. The key idea in this work is that initiating from a more informative source than Gaussian enables diffusion methods to mitigate the above limitations. We contribute both theoretical results, a new method, and empirical findings that show the benefits of using an informative source policy. Our method, which we call BRIDGeR, leverages the stochastic interpolants framework to bridge arbitrary policies, thus enabling a flexible approach towards imitation learning. It generalizes prior work in that standard Gaussians can still be applied, but other source policies can be used if available. In experiments on challenging simulation benchmarks and on real robots, BRIDGeR outperforms state-of-the-art diffusion policies. We provide further analysis on design considerations when applying BRIDGeR.
@@ I Introduction
An examination of existing diffusion-style imitation learning reveals a fundamental issue: these models learn to shape a policy starting from standard Gaussian noise, which is often starkly different from the intended policy or action distribution. The key insight in our work is that initiating from Gaussian noise isn’t a prerequisite. To explore this, we move beyond the conventional diffusion framework and employ stochastic interpolants [2] for bridging arbitrary densities within finite time (Fig. 1). This approach allows us to leverage stochastic source policies, enabling the diffusion process to begin from a more informative starting point. Example source policies include policies hand-crafted using prior knowledge of the task or data-driven policies trained on a similar task. We find that this shift retains the inherent advantages of diffusion-style imitation learning, but positively impacts inference time and performance. Practically, this leads to faster generation and more accurate robot actions. Our approach also generalizes prior work in diffusion-based imitation learning since if no source policy is available, simple distributions such as the Gaussian can be used.
@@
In this paper, we first contribute a theoretical analysis of the impact of different source policies in diffusion. In brief, we find that under reasonable assumptions, selecting a better source policy results in better target policies. We then turn to a practical approach for incorporating source policies into diffusion methods. Applying the stochastic interpolants framework [2] to imitation learning, we derive a new method called BRIDGeR (Behavioral Refinement via Interpolant-based Diffusion for Generative Robotics). To our knowledge, our work is the first adaptation of this bridging methodology to imitation learning, contrasting with its previous use in simple synthetic tasks [60] and image generation [60, 25]. In addition to standard neural architecture design for the learnt forward model, the stochastic interpolant framework relies on several critical design choices including the source policy and interpolant. The interpolant dictates how a sampled point transitions from the source to the target distribution, with the transition modulated by noise introduced through time-dependent Gaussian latent variables [2]. Intuitively, the interpolant forms a “bridge” or “guide” between two policies (e.g., it gradually changes poor robot actions into better ones).
@@
We contribute a systematic empirical study of the effects of using source policies (and other design elements) on a diverse set of robot tasks, including the Franka kitchen benchmark, grasp generation, and manipulation using a robot hand. Overall, the experimental results coincide with our theoretical findings; Gaussians were seldom the most effective source distribution and surprisingly, even simple heuristic distributions resulted in superior learnt policies compared to the Gaussian. We demonstrate that given a good source policy, BRIDGeR surpasses existing state-of-the-art diffusion policies. Additionally, we discuss the effects of the interpolant function when learning highly multi-modal behaviors. Similar positive results were observed on real-world experiments using two robots: a Franka Emika Panda arm with a two-finger gripper for stable grasping, and a UR5e equipped with a Shadow Dexterous Hand Lite for synthetic wound cleaning. These tasks involved real-world high-dimensional observations (e.g., point clouds and images) and complex actions (22 action dimensions per time-step for the wound cleaning task).
@@
II Preliminaries: Background & Related Work
@@ II-A Problem Formulation
In imitation learning [20, 59, 51], we wish to learn a policy from expert demonstrations. Let $\pi_{1}(a|x)$ be an expert policy; it captures the probability of an expert selecting an action $a$ given an observation $x$. Suppose we have a dataset $\mathcal{D}=\{\smash{x^{(i)}},\smash{a^{(i)}}\}_{i=1}^{N}$ drawn from $\pi_{1}(a|x)$ and let $\pi_{0}(a|x)$ be a distribution that we can easily sample $a$ from.
@@ II-B Diffusion-based Policy Learning
II-C Related Work
BRIDGeR builds upon a large body of work in imitation learning, specifically behavior cloning. Behavior cloning has had a long history, with early methods adopting supervised learning techniques, principally regression [5, 34, 34, 58, 13, 38, 48]. Although efficient, regression-based methods were generally unable to capture multi-modal behavior. Later methods used classification methods, by discretizing the action space [41, 42, 4, 52], to address this limitation but were sensitive to hyperparameters (such as the level of discretization) and struggled with high-precision tasks [12].
Modern imitation methods employ generative models, starting with Gaussian Mixture Models [28], then transitioning to more powerful Energy-Based Models (EBMs) [8, 12, 21, 44] and Diffusion Models [40, 7, 32, 31]. Compared to EBMs, diffusion models have demonstrated better training stability [7] and effectiveness in generating consistent, multi-modal action sequences for visual-motor control [40, 7, 32, 31] and hierarchical, long-horizon planning tasks [36, 16, 30, 6, 53]. Recent progress has been made in reducing the computational costs of diffusion methods by adapting the number of diffusion steps [18] or leveraging DDIM [37].
Unlike the above methods, BRIDGeR is the first to generalize diffusion-type policy learning to exploit source stochastic policies. As we will see, this leads to better performance compared to strong baselines (such as DDIM) when informative policies are available, especially when using a small number of diffusion steps. Since BRIDGeR adapts existing distributions, it can be seen as a few-shot learner, specifically for imitation learning [10, 51]. Existing work on few-shot imitation learning requires either specific prior policies [11, 57, 24] or a hierarchical structure [54]. In contrast, BRIDGeR only requires that we are able to sample from the source policy. For fair comparison against few-shot methods, we adopt residual learning which is used for few-shot reinforcement learning [43, 1, 22], planning [29], and shared autonomy [39, 56]. BRIDGeR is also related to recent gradient-flow methods [3, 47] but the SDE is over finite time and does not use classifiers to approximate the drift.
@@ IV-A Stochastic Interpolants for Imitation Learning
Let $(C^{r}(\mathbb{R}^{n}))^{m}$ be the space of $r$ continuously differentiable functions from $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$ and $(C_{0}^{r}(\mathbb{R}^{n}))^{m}$ as the space of compactly supported and $r$ continuously differentiable functions from $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$. We extend the definition of stochastic interpolant in [2] to condition on observation $x$ (e.g., an image or joint angles), similar to the concurrent work [19].
@@
     $L_{b}(\theta)=\frac{1}{N}\sum\left(b_{\theta}(t,a_{t},x)-(\partial_{t}I+\dot{\gamma}z)\right)^{2}$
     $L_{s}(\eta)=\frac{1}{N}\sum\left(s_{\eta}(t,a_{t},x)+\gamma^{-1}z\right)^{2}$
     $L_{v}(\phi)=\frac{1}{N}\sum\left(v_{\phi}(t,a_{t},x)-\partial_{t}I\right)^{2}$
@@ IV-B Design Decisions
V Experiments
This section describes experiments designed to evaluate the performance of BRIDGeR relative to recent methods, particularly diffusion-based imitation learning. More importantly, we aimed to test our hypothesis that leveraging better source distributions within a diffusion framework leads to better policies. We further hypothesized that BRIDGeR performs better than DDIM policies given a small number of diffusion steps, and limited data. Finally, we sought to examine the effect of decision decisions, particularly the interpolant choice. We first give an overview of our experimental setup (details relegated to Appendix) followed by our main results.
V-A Domains
@@
Table I: Average task performance on Adroit (success rate) and Franka Kitchen (number of successful sub-tasks). Best scores in bold. We compare BRIDGeR against state-of-the-art methods under a different number of diffusion steps when trained with the Large dataset. BRIDGeR with $k=0$ indicates the source policy. BRIDGeR generally outperforms the competing methods. Results are similar for the small and medium datasets with complete results in the Appendix.
@@
Table II: Success rate (averaged over $100$ grasps on ten test objects). BRIDGeR significantly outperforms DDIM and Residual Policy across the number of diffusion steps. Compared to SE3, BRIDGeR achieve higher success rate when the number of diffusion steps is small. We show up to $k=160$ steps to be consistent with prior reported results [49]. Best scores in bold.
@@
V-D Main Results and Discussion
Figure 7: Fifteen sampled grasps across the diffusion steps for BRIDGeR, DDIM and SE3. We visualize $15$ grasps samples every two diffusion steps until $k=10$.
@@ V-D Main Results and Discussion
Figure 9: Average Earth Mover’s Distance between generated grasps pose and target grasp poses on Seen Categories in Grasp Generation task ($k=160$ for SE3, DDIM and BRIDGeR). Lower EMDs indicate BRIDGeR better mimics the dataset distribution.
Figure 10: Task performance using different source policies. Each colored line shows the performance of a source policy before (“Source”) and after diffusion (“Generated Target”).
@@
VI Real World Robot Experiments
In this section, we present findings from experiments aimed at evaluating BRIDGeR in real-world domains with noisy and high-dimensional observations (point-cloud and image). As in the previous section, we convey our main results and refer readers to the appendix for further details. Given our simulation results, we hypothesized that BRIDGeR would outperform existing state-of-the-art diffusion-based methods under a computation budget and conducted experiments using two tasks:
Figure 13: (A) Real-world Grasping using a Panda arm with a two-finger gripper. Observations were point clouds obtained from the RealSense Camera on the robot (B) Test objects used in our experiments (unseen during training). (C) Grasp samples from the competing models on three objects (20 diffusion steps).
Table III: Real-World Grasping Success rate (averaged over 10 grasps on 10 test objects). Best scores shown in bold.
BRIDGeR SE3
CVAE Heuristic
$k=0$ $0.07\pm 0.12$ $\mathbf{0.02\pm 0.00}$ $00.00\pm 0$ 0.00
$k=5$ $0.59\pm 0.17$ $\mathbf{0.66\pm 0.13}$ $0.05\pm 0.07$
$k=20$ $0.71\pm 0.18$ $\mathbf{0.79\pm 0.18}$ $0.56\pm 0.20$
$k=160$ $0.75\pm 0.18$ $\mathbf{0.81\pm 0.24}$ $0.73\pm 0.21$
6-DoF Grasping (Point Cloud Observations) where a Franka-Emika Panda arm has to grasp and lift objects (Fig. 13). The experiment involved $10$ everyday objects with 10 grasp trials per object. Objects were perceived using a RealSense camera that provided point cloud observations. We used the models obtained from our simulation experiments and compared BRIDGeR (with the Power3 interpolant) against the SE3 diffusion model. We used each method to generate a set of end-effector grasp poses, and the MoveIt motion planner to plan and execute a collision-free path to grasp and lift the objects. A trial was considered successful if the robot managed to lift the object without it falling out of the robot’s grasp.
Cleaning (Image and State Observations). Inspired by assistive tasks in healthcare such as wound cleaning and bed-bathing, we used a Shadow Dexterous Hand Lite mounted on a UR-5e arm to perform a synthetic wound cleaning task. The robot had to grasp a sponge and wipe marks off a surgical practice model that mimics human tissue (Fig. 14). This task is challenging as it involves complex high-dimensional action sequences (22 dimensions per time-step, 48 time-steps per prediction), multi-modal demonstrations (e.g., wiping off a mark either from left to right, or from top to bottom), and manipulating a deformable sponge. The mark can be at one of nine possible locations and the robot has to learn appropriate behavior conditioned on visual observations (a RGB image from a RealSense Camera) and its current joint angles. We compared BRIDGeR against Diffusion Policy [7] (DDIM) using a normalized cleaned area score, which represents how much of the mark was wiped off. We apply receding-horizon control; the models predict $48$ action steps, of which $16$ steps of actions are executed on the robot without re-planning. The models were trained using 60 demonstrations provided via kinesthetic teaching and replay.
Figure 14: (A) Real-world Synthetic Wound Cleaning using a UR5e with Shadow Dextrous Hand Lite. (B) Demonstrations consisted of moving the hand to the sponge from an initial position, grasping it, then manipulating the sponge to wipe off the mark. Initial positions were randomized in a region roughly 20-30cm above the sponge and the mark was in one of 9 possible positions. The robot had to learn an action policy conditioned upon RGB images from the RealSense Camera and its joint angles (both arm and hand).
Table IV: Normalized cleaned area (averaged over 9 positions) for the Cleaning Task. Best scores shown in bold.
BRIDGeR DDIM
CVAE Heuristic
$k=0$ $00.00\pm 0$ 0.00 $\mathbf{0.14\pm 0.31}$ $00.00\pm 0$ 0.00
$k=5$ $0.46\pm 0.42$ $\mathbf{0.50\pm 0.35}$ $0.19\pm 0.25$
$k=20$ $0.47\pm 0.39$ $\mathbf{0.51\pm 0.33}$ $0.27\pm 0.33$
$k=80$ $\mathbf{0.32\pm 0.31}$ $0.31\pm 0.29$ $0.16\pm 0.15$
Table V: Average time to successfully grasp the sponge and roughness of generated action sequences.
BRIDGeR DDIM
CVAE Heuristic
Time (Seconds) $14.25\pm 3.09$ $\mathbf{13.00\pm 1.87}$ $36.25\pm 11.17$
Roughness $0.12\pm 0.01$ $\mathbf{0.08\pm 0.01}$ $0.34\pm 0.02$
VI-A Results
BRIDGeR outperforms the baselines, with larger gaps when the number of diffusion steps is small. For the grasping task, BRIDGeR achieves significantly higher success rates compared to SE3 (Table. III). Qualitatively, we observed the grasps generated by BRIDGeR to be more accurately positioned (samples shown in Fig. 13), which led to more stable grasps.
For the Cleaning task, Table. IV shows that BRIDGeR was better at wiping off the marks compared to the DDIM diffusion policy. Interestingly, performance for all models fell at the largest number of diffusion steps ($k=80$); a potential cause is that errors accumulate during diffusion given the high-dimensional actions. Nevertheless, BRIDGeR attains significantly higher scores. Qualitatively, we observed BRIDGeR generates smoother action trajectories (please see the accompanying supplementary videos for examples). At $k=5$, the DDIM model produced jerkier behavior, with random movements in the arm and fingers. In contrast, BRIDGeR’s trajectories better mimicked the demonstrations, which led to faster completion of the task. Table V summarizes this observation quantitatively; we see BRIDGeR was quicker to grasp the sponge and had lower trajectory roughness (the average norm of the second derivative of the action trajectories).
@@ VII Conclusions and Future Work
In this work, we investigate the potential of integrating informative source distributions into diffusion-style imitation learning. We provide theoretical results that support this idea and propose BRIDGeR, a stochastic interpolant method for imitation learning. Our experiments results show that BRIDGeR outperforms strong baselines, including state-of-the-art diffusion policies on various benchmark tasks and real-world robot experiments. We provide additional analyses on our experimental results to elucidate the effect of various design decisions within BRIDGeR.
Limitations and Future Work. Here, we have shown that leveraging prior knowledge (in the form of source policies) improves learned diffusion policy performance. BRIDGeR opens up avenues to explore different forms of prior knowledge, e.g., policies designed for other tasks (transfer learning) and those constructed using foundation models. Specific to BRIDGeR, we explored a salient but limited set of design considerations; future work can examine more elaborate interpolant functions, along with more in-depth analysis of noise schedules and diffusion coefficients. Finally, we plan to extend BRIDGeR incorporate considerations such as safety constraints.